two integers differ by 1 the sum of their reciprocals equals to 7 divided by their product.find the two integers.
Answer by mananth(16949)

x-y = 1................1
1/x +1/y = 7/xy
(x+y)/xy = 7/xy
x+y =7................2
add (1) & (2)
2x=8
x=4
y=3